OpenStudy (anonymous):

please help MEDALS WILL BE GIVEN!! A . d <7 B . <-0.9 C . d > 7 D . d > -1

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }d-4 \frac{ 1 }{ 5 } > 3\]

OpenStudy (anonymous):

@Giggglez

OpenStudy (anonymous):

uhh? I dont know hun lol sorry..

OpenStudy (anonymous):

it fine

OpenStudy (anonymous):

@Aeon_Valentine00 @EvonHowell @Hero @iJustAteCheese @itsmylife @oldrin.bataku @poopsiedoodle @thivitaa @x1xofxaxkndx

OpenStudy (anonymous):

@abb0t

OpenStudy (nottim):

@NotTim

OpenStudy (anonymous):

\[4\frac{ 1 }{ 5 }=\frac{ 21 }{ 5 }\]\[\frac{ 4 }{ 5 }d-\frac{ 21 }{ 5}>3 \]

OpenStudy (anonymous):

Don't mass tag...

OpenStudy (anonymous):

now you can do that

OpenStudy (anonymous):

i dont understand how

OpenStudy (anonymous):

\[4\frac{ 1 }{ 5 }=\frac{ 4*5+1 }{ 5 }=\frac{ 20+1 }{ 5 }=\frac{ 21 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }d>\frac{ 21 }{ 5 }+3\]

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

please help

OpenStudy (nottim):

maybe ask what you are having problems with this question specificly.

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

hmm good point @NotTim

OpenStudy (anonymous):